### v1

# Persistent Cost of Lipschitz Maps

Abstract
A 1-Lipschitz map between compact metric spaces $f:X\to Y$ induces a homomorphism on the level of persistent d-homology of the associated Vietoris-Rips simplicial filtrations. This homomorphism naturally bounds the Bottleneck distance between said modules in terms of the maximum size of its kernel and cokernel persistence modules, a stable invariant which we refer to as the persistent cost associated to $f$. We provide metric upper bounds on the persistent cost of a 1-Lipschitz map and revisit its stability.
Introduction
Persistent homology is a fundamental tool in the emerging field of topological data analysis (TDA) that extracts a vast amount of metric information by encoding homological features across scales. The richness of this output coupled with its explainability has yielded significant insights across a spectrum of disciplines, including biology and time series analysis, among others, c.f.[otter2017roadmap, perea2015sliding, gardner2022toroidal], and been integrated into learning pipelines, c.f., [hensel2021survey, hofer2017deep].
We outline the standard pipeline for topological data analysis. Given a metric space $X$, the Vietoris–Rips filtration $(\operatorname{VR}_{s}(X))_{s\geq 0}$ is constructed by including simplices whose vertex set has diameter at at most $s\geq 0$. As the parameter increases, the complex evolves from a discrete set to the full power set of $X$, encoding geometric information at multiple scales. Applying the homology functor in a fixed dimension $d$ yields a persistence module: a sequence of vector spaces connected by linear maps that track the birth, persistence and death of homological features.
While the TDA community has primarily focused on analyzing individual point clouds, the functorial nature of these constructions naturally leads us to consider mappings between datasets. Indeed, we advocate for a treatment of maps between spaces that parallels that of the spaces themselves, enabling us to address the geometry of functions from a persistent homology perspective. In this context we shall consider the natural case of non-expanding, i.e., 1-Lipschitz, maps $f:X\to Y$ between Vietoris-Rips filtrations of finite metric spaces which yields a homomorphism of point-wise finite dimensional persistence modules,
$f_{*}:\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}}(X))\to\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}}(Y)).$
A homomorphism of persistence modules defines kernel, cokernel, and image modules via point-wise linear algebra, along with naturally induced maps. The “size” of the modules $\operatorname{Ker}(F)$ and $\operatorname{Coker}(F)$, measured by their interleaving distance to the trivial module $\Delta$, serves as a measure of how far the map $F$ is from a persistent isomorphism. This leads us to define the persistent cost of a $1$-Lipschitz function as the maximum interleaving distance of kernel and cokernel persistence modules to the diagonal,
$\begin{array}[]{rcl}C(f)&=&\max\big\{d_{I}(\operatorname{Ker}(f),\Delta)\penalty 10000\ ,\penalty 10000\ d_{I}(\operatorname{Coker}(f),\Delta)\big\}\end{array}$ (1)
This persistent cost bounds the interleaving distance between the modules associated to $X$ and $Y$, by the Induced Matching Theorem of U. Bauer and M. Lesnick in [bauer2015induced] (see also [bauer2020persistence]).
Building on this result, we contribute a metric upper bound for the persistent cost of non-expanding map $f:X\to Y$, showing that $C(f)$ is bounded above by its distortion and the Hausdorff distance between its image and the codomain, $f(X)\subseteq Y$. Specifically, we establish the second inequality in the following sequence.
$d_{I}\left(\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}}(X)),\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}}(Y))\right)\penalty 10000\ \leq\penalty 10000\ C(f)\penalty 10000\ \leq\penalty 10000\ \operatorname{dist}(f)+2\cdot d_{H}(f(X),Y)$ (2)
From a purely algebraic perspective, as we range over all possible persistence module homomorphisms $F_{\bullet}:V_{\bullet}\to W_{\bullet}$, the infimum of the associated persistent cost attains the interleaving distance between the given modules. However, our interest lies in geometrically defined persistence homomorphisms, i.e. those induced by $1$-Lipschitz maps between the underlying spaces via the homology of their Vietoris-Rips filtrations. In this geometric context, we provide an example of two finite metric spaces that yield identical persistence diagrams while any 1-Lipschitz map is bounded away from zero in its associated persistent cost.
A prior concern is the stability of kernel and cokernel persistence modules which was addressed in [cohen2009persistent]. We provide a self-contained exposition of this fact that was claimed without proof. In doing so, we introduce a Gromov-Hausdorff type distance between maps $f:X\to Y$, $f^{\prime}:X^{\prime}\to Y^{\prime}$, of compact metric spaces that may be of independent interest.
Our results are developed within the framework of point-wise finite dimensional persistence modules and finite metric spaces, and naturally extended to q-tame modules that arrise from compact metric spaces.
The paper is organized as follows. A first section provides a self-contained exposition of the algebraic notions regarding persistence theory. We then move on to the geometric of spaces, maps and distances. The last two subsection are focused on our contributed results on metric upper bounds for the cost, Theorem 2.11 and stability, Theorem 2.16.
1 Algebraic Persistence
The main references for this section are the comprehensive book by S. Oudot in [oudot2015persistence] and the papers on the Induced Matching Theorem [bauer2015induced, bauer2020persistence] by U. Bauer and M. Lesnick.
1.1 Persistence Modules
A persistence module is a linear representation of $\mathbb{R}$, or a given subset $T\subseteq\mathbb{R}$, regarded as a poset. In other words, it is a functor
$V:(\mathbb{R},\leq)\to\mathrm{Vect}_{\mathbb{F}},$ (3)
where $\mathrm{Vect}_{\mathbb{F}}$ denotes the category of vector spaces over a field $\mathbb{F}$ with linear transformations as homomorphisms.
This, a persistence module is more than just a one-parameter family of vector spaces; as a representation of a poset, it carries “inner persistence” or “transition” maps
$\varphi_{s<t}:V_{s}\xrightarrow{\penalty 10000\ \penalty 10000\ }V_{t},$
for each pair $s<t\in\mathbb{R}$, which satisfy
$\varphi\circ\varphi=\varphi,$
the last equation being read as $\varphi_{s_{2}<s_{3}}\circ\varphi_{s_{1}<s_{2}}=\varphi_{s_{1}<s_{3}}$.
As a first and computationally relevant approach to the subject, the reader may consider the case of finite-dimensional spaces indexed over a finite set $0=\epsilon_{0}<\epsilon_{2}<\dots<\epsilon_{r}$, giving rise to
$V_{\epsilon_{0}}\to V_{\epsilon_{1}}\to\cdots\to V_{\epsilon_{r}}.$ (4)
The key idea of persistence is that the dimension of $V_{s}$ may be preserved (via an isomorphism) in $V_{t}$ for $t>s$.
However, indexing over the real numbers has several advantages. It is less cumbersome, as it avoids the problem of refining different indexing sets. It is also more general, since any finitely indexed persistence module can be easily extended to a $\mathbb{R}$-indexed one.
We will first develop the theory within the setting of point-wise finite dimensional (p.f.d.) modules indexed over $\mathbb{R}$, as in definition (3). A fundamental result states that such persistence modules admit a unique decomposition into interval modules [crawley2015decomposition, oudot2015persistence]. Specifically,
$V\cong\bigoplus_{I\in\mathcal{I}}\mathbb{F}^{I},$ (5)
where $\mathcal{I}$ is a collection of intervals $I\subseteq\mathbb{R}$ each defining an “interval module” as follows:
$(\mathbb{F}^{I})_{(t)}=\begin{cases}\mathbb{F},&t\in I,\\ 0,&\text{otherwise},\end{cases}$
with inner persistence maps:
$\varphi_{s<t}=\begin{cases}Id,&s,t\in I,\\ 0,&\text{otherwise}.\end{cases}$
The field $\mathbb{F}$ will play no role in the discussion, usually being $\mathbb{Z}_{p}$ for p prime, $\mathbb{Q}$ or $\mathbb{R}$ in practice.
The multiset of intervals $\mathcal{I}$ is called the persistence diagram (or barcode). Each interval $\mathcal{I}$ is represented buy its endpoints $(b_{i},d_{i})\in\bar{\mathbb{R}}^{2}$. A barcode represents each interval as a horizontal line from $b$ to $d$, as shown in Figure 1.
bbdd
12345tt
Figure 1: A persistence diagram and barcode representation of the same decomposition.
In order to define a distance between persistence diagrams $D_{1},D_{2}$, we may begin by setting a partial matching between their persistent features. If a pair $(b_{1},d_{1})\in D_{1}$ and $(b_{2},d_{2})\in D_{2}$ are matched, we may ponder their dissimilarity by the quantity $(b_{1}-b_{2},d_{1}-d_{2})||_{\infty}$. Unmatched elements $(b,d)\in D_{i}$ are virtually paired to an ephemerous feature $(\frac{b+d}{2},\frac{b+d}{2})$, i.e., to an element listed in neither diagram, thus weighting in as $\frac{(d-b)}{2}$. In other words, for a partial matching $\Theta$ between diagrams $D_{1},D_{2}$ we have defined its dissimilarity as
$diss(\Theta_{D_{1},D_{2}})=\max\Bigg\{\underset{(b_{1},d_{1})\sim_{\Theta}(b_{2},d_{2})}{\sup\quad\quad}||(b_{1}-b_{2},d_{1}-d_{2})||_{\infty}\penalty 10000\ \penalty 10000\ ,\penalty 10000\ \penalty 10000\ \penalty 10000\ \underset{(b,d)\notin\pi_{i}(\Theta)}{\sup\quad\quad}\frac{(d-b)}{2},\penalty 10000\ \Bigg\}$
Definition 1.1.
The Bottleneck distance between persistence diagrams $D_{1},D_{2}$ is defined as the infimum dissimilarity over all partial matchings, i.e.,
$d_{Bottl}(D_{1},D_{2})=\underset{\theta}{\inf}\penalty 10000\ diss(\theta_{D_{1},D_{2}})$ (6)
Notice that any interval decomposable persistence module is at zero distance of a module constructed using only intervals of the form $[a,b)$. We will comment on this in Section 1.4
1.2 Homomorphisms and Interleavings
Definition 1.2.
A homomorphism $F:V_{\bullet}\to W_{\bullet}$ between persistence modules is a natural transformations between $V_{\bullet}$ and $W_{\bullet}$ regarded as functions. In other words, a homomorphism is a collection of linear maps $F_{\bullet}$, at each scale, commuting with inner maps of each module, so that for $s<t$ we have:
${V_{s}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{V}_{s<t}}$$\scriptstyle{F_{s}}$${V_{t}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F_{t}}$${W_{s}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{W}_{s<t}}$${W_{t}.}$
In short, a persistence homomorphism.
For $\epsilon\in\mathbb{R}$, a $\epsilon$-degree homomorphism between given modules $V_{\bullet}$ and $W_{\bullet}$, is a standard persistence module homomorphism when the codomain is shifted by $\epsilon$, namely $F:V_{\bullet}\to W_{{}_{\bullet}+\epsilon}$.
${V_{s}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{V}}$$\scriptstyle{F_{s}}$${V_{t}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{F_{t}}$${W_{s+\epsilon}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{W}}$${W_{t+\epsilon}.}$
In practice, it is $\epsilon\geq 0$, though it is not necessary.
We may compare modules by means of interleavings which are defined as follows.
Definition 1.3.
A $\delta$-interleaving between persistence modules $V_{{}_{\bullet}}$ and $W_{{}_{\bullet}}$ is a pair of degree $\delta>0$ homomorphisms $F:V_{\bullet}\to W_{\bullet}$, $G:W_{\bullet}\to V_{\bullet}$, such that the following diagrams commute:
[…]
Whenever clear from context we omit indexing homomorphisms and naming inner persistence maps of each module.
Definition 1.4.
The interleaving distance between persistence modules $V_{\bullet},W_{\bullet}$ is defined by:
$d_{I}(V_{{}_{\bullet}},W_{{}_{\bullet}})=\inf\{\epsilon:\text{there is an $\epsilon$-interleaving between $V_{\bullet}$ and $W_{\bullet}$}\}$
The interleaving distance agrees with the Bottleneck distance between associated diagrams, in the interval decomposable case, i.e.,
$d_{Bottl}(diag(V_{\bullet}),diag(W_{\bullet})=d_{I}(V_{\bullet},W_{\bullet}).$
This result is known as the algebraic isometry theorem, see [lesnick2015theory, oudot2015persistence].
1.3 Kernel, cokernel and the Induced Matching Theorem
Given a homomorphism $F:V_{\bullet}\to W_{\bullet}$, we can define persistence modules for the point-wise kernel, cokernel, and image. Standard diagram chasing shows that these form an exact sequence of persistence modules.
$0\to\operatorname{Ker}(F)_{\bullet}\xrightarrow{\penalty 10000\ \penalty 10000\ \penalty 10000\ }V_{\bullet}\xrightarrow{\penalty 10000\ \penalty 10000\ F\penalty 10000\ \penalty 10000\ }W_{\bullet}\xrightarrow{\penalty 10000\ \penalty 10000\ \penalty 10000\ }\operatorname{Coker}(F)_{\bullet}\to 0.$ (7)
The following result describes how a homomorphism induces a matching between the interval decompositions of $V_{\bullet}$ and $W_{\bullet}$.
Theorem 1.5 (Induced Matching Theorem, [bauer2015induced, bauer2020persistence]).
Let $V_{\bullet},W_{\bullet}$ be p.f.d. persistence modules and $F_{\bullet}:V_{\bullet}\to W_{\bullet}$ a homomorphism. Then there exists an induced matching between intervals in their respective interval decompositions satisfying the following:
  1. 1.
    a matched pair of bars $[a,b),[c,d)$ in the respective decompositions of $V_{\bullet}$ and $W_{\bullet}$, satisfies $a\leq c<b\leq d$;
  2. 2.
    the unmatched intervals in $V_{\bullet}$, as well as the non-overlapping segment $[a,c)$ of a pair of matched bars, correspond both to intervals in $\operatorname{Ker}_{\bullet}$/
  3. 3.
    analogously, the unmatched intervals in $W_{\bullet}$, as well as the non-overlapping segment $[b,d)$ of a pair of matched bars, correspond both to intervals in $\operatorname{Coker}_{\bullet}$;
  4. 4.
    the middle portion $[c,b)$ in common to a matched pair of bars.
1.4 The right setup
We conclude this algebraic section by reviewing the framework of q-tame persistence modules, following [chazal2014observable] and [oudot2015persistence].
A q-tame persistence module over $\mathbb{R}$ is a slightly more general notion of persistence module where the point-wise finite dimensional hypothesis is replaced by the weaker requirement that every inner persistence map $\varphi_{s<t}$ be of finite rank.
A p.f.d. module can be retrieved from a q-tame one, for example one could consider $\hat{V}_{t_{0}}:=\lim_{s\to{t_{0}}^{+}}\operatorname{Im}(\varphi_{t_{0}<s})$. In fact there is a quotient category, the observable category, retrieved from the stabilization of persistence modules up to direct sum with the “ephemeral” subcategory given by those persistence modules all of whose inner homomorphisms are trivial, i.e.,
$\mathbf{Obs}:=\mathbf{Pers}/\mathbf{Eph},$
where $\mathbf{Pers}:=[(\mathbb{R},\leq),\mathbf{Vect}_{\mathbb{F}}]$ as before. Ephemerous modules correspond to “still-born” persistence features of the form $[a,a]$.
Every q-tame persistence module is isomorphic to a p.f.d. one in $\mathbf{Obs}$, particularly to a unique interval-decomposable one with intervals of chosen decoration, i.e., type of strict or relaxed condition on endpoints: $[a,b],(a,b),[a,b)$ or $(a,b]$. The barcode associated to a q-tame module is understood to be that of an interval decomposition of its associated p.f.d. observable.
The homology of the Vietoris-Rips filtration over compact metric may not yield a p.f.d. persistence module but a q-tame one nonetheless.
Throughout the text "distance" is understood in the general sense as a symmetric non-negative extended real function on pairs for which the triangular inequality holds. However, it may only be a pseudo-distance in the sense that there is a degenerate non-unique set of points at zero distance. Considering the observable category solves the issue, as the pseudo-distance becomes a true distance once zero interleaved p.f.d. persistence modules become isomorphic in $\mathbf{Obs}$, see [lesnick2015theory]. In particular, this makes the already introduced graphic representation as persistence diagrams or barcodes a faithful one, as in previous Figure 1.
2 Persistent cost
A general reference for this section is Oudot’s book [oudot2015persistence] with regards to persistence theory and the references therein. For metric arguments we refer to [burago2001metricgeometry].
2.1 Vietoris-Rips filtration and maps
The Vietoris-Rips construction is a standard procedure by which a filtration of simplicial complexes is produced from an actual dataset, presented as a finite metric space. To be precise, for each $\epsilon\geq 0$ define:
$\mathrm{VR}_{\epsilon}(X)=\{\sigma\subseteq X\mid d(x,y)\leq\epsilon,\,\forall x,y\in\sigma\}\subseteq\mathcal{P}(X),$
i.e., a collection of subsets of $X$. Applying the d-dimensional homology functor $H_{d}(-,\mathbb{F)}$, with coefficients in a field $\mathbb{F}$, yields a persistence module $V_{\bullet}:=\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))$ with inner persistence maps $\varphi_{s<t}:=(\iota_{s<t})_{*}$.
Let us consider perturbations of $1$-Lipschitz functions as follows.
Definition 2.1.
A function between metric spaces, $h:X\to Y$, is said to be an $\epsilon$-quasi $1$-Lipschitz map if it satisfies the following inequality:
$d_{Y}(h(x),h(x^{\prime})\leq d_{X}(x,x^{\prime})+\epsilon.$
The Vietoris-Rips construction is particularly well suited for our metric endeavors, as follows from the following.
Lemma 2.2.
Let $h:X\to Y$ be an $\epsilon$-quasi $1$-Lipschitz map. Then $h$ induces an $\epsilon$-degree persistence homomorphism: $h_{*}:H_{d}(\operatorname{VR}_{{}_{\bullet}}(X))\to H_{d}(\operatorname{VR}_{{}_{\bullet}+\epsilon}(Y)).$
Moreover, if two $\epsilon$-quasi $1$-Lipschitz maps $h_{1},h_{2}:X\to Y$ are within uniform distance $\delta\geq 0$, then they induce the same $(\epsilon+\delta)$-degree map on homology.
Proof.
The map $h$ induces a simplicial map:
$\hat{h}:\operatorname{VR}_{{}_{\bullet}}(X)\to\operatorname{VR}_{{}_{\bullet},+\epsilon}(Y),\quad[x_{1}\dots x_{r}]\mapsto[h(x_{1})\dots h(x_{r})],$
since a set $\sigma\subseteq X$ of diameter $\ell$ maps to a set $h(\sigma)\subseteq Y$ of diameter at most $\ell+\epsilon$. This commutes with inclusions and thus gives a morphism between filtrations. The homology functor yields the desired persistence homomorphism.
For the second claim, suppose $d_{Y}(h_{1}(x),h_{2}(x))\leq\delta$ for all $x\in X$. For any simplex $\sigma\in\operatorname{VR}_{{}_{\bullet}}(X)$ of diameter $\ell$, consider the union:
$h_{1}(\sigma)\cup h_{2}(\sigma)\subseteq Y.$
For any $y,y^{\prime}\in h_{1}(\sigma)\cup h_{2}(\sigma)$, we have: - If both come from the same $h_{i}$, then $d(y,y^{\prime})\leq\ell+\epsilon$. - If $y=h_{1}(x)$, $y^{\prime}=h_{2}(x^{\prime})$, then:
$d(y,y^{\prime})\leq d(h_{1}(x),h_{2}(x))+d(h_{2}(x),h_{2}(x^{\prime}))\leq\delta+(\ell+\epsilon).$
Thus, the diameter of $h_{1}(\sigma)\cup h_{2}(\sigma)$ is at most $\ell+\epsilon+\delta$. This implies that $\hat{h}_{1}$ and $\hat{h}_{2}$ are contiguous as maps:
$\operatorname{VR}_{{}_{\bullet}}(X)\to\operatorname{VR}_{{}_{\bullet}+\epsilon+\delta}(Y),$
and therefore induce the same map on homology. ∎
Remark 2.3.
If $f:X\to Y$ is a $1$-Lipschitz map, it induces a persistence homomorphism:
$f_{*}:H_{d}(\operatorname{VR}_{{}_{\bullet}}(X))\to H_{d}(\operatorname{VR}_{{}_{\bullet}}(Y)).$
We may then define the persistent cost of $f$ as:
$C(f):=C(f_{*})=\max\left\{d_{I}(\operatorname{Ker}(f_{*}),\Delta),\ d_{I}(\operatorname{Coker}(f_{*}),\Delta)\right\},$
where $\Delta$ is the trivial persistence module.
As natural examples one should consider the isometric inclusion of a sample withing a larger space $S\subseteq Y$, a clustering assignment $p:X\to X/\sim$ with a convenient metric or, in fact, any map $f:X\to Y$ between finite metric spaces as long as the distance of the target space $Y$ is conformally adjusted to ensure that the 1-Lipschitz condition is satisfied.
A source of $\epsilon$-quasi $1$-Lipschitz maps with non-trivial $\epsilon>0$ is given by the choice of functions subordinated to a given correspondence between metric spaces.
Definition 2.4.
A correspondence between two sets is a relation between two subsets, $\mathcal{C}\subseteq X\times Y$, that in fact relates every element of the factors, i.e., such that $\pi_{X}(\mathcal{C})=X$, $\pi_{Y}(\mathcal{C})=Y$.
When these spaces $x,y$, are metric we have a notion of distortion given by the supremum of the difference between the distances of related pairs, i.e.,
$dist(\mathcal{C})=sup\Bigg\{|d(x,x^{\prime})-d(y,y^{\prime})|\penalty 10000\ :\penalty 10000\ (x,y),(x^{\prime},y^{\prime})\in\mathcal{C}\Bigg\}$ (8)
Within this framework we may state the Gromov-Hausdorff classical stability and formulate the proof of our main theorem.
Definition 2.5.
Given metric spaces $X$, $Y$, consider all possible correspondences and take half the infimum over their distortions to be the Gromov-Hausdorff distance between them,
$d_{GH}(X,Y)=\frac{1}{2}\,\underset{\mathcal{C}}{\inf}\{\!\penalty 10000\ dist(\mathcal{C})\!\penalty 10000\ \}.$
The alternative, albeit original, definition is to consider the infimum over all metric spaces $Z$ and isometric embeddings $f:X\to Z$, $g:Y\to Z$, of the quantity $d_{\mathrm{H}}^{Z}$ given by the Hausdorff distance in $Z$ of the respective images from $X,Y$.
$d_{\mathrm{GH}}(X,Y)=\inf_{Z,f,g}d_{\mathrm{H}}^{Z}(f(X),g(Y)).$
The Gromov-Hausdorff distance is a distance when we consider compact subsets up to isometry.
Corollary 2.6.
A correspondence with $\epsilon$-distortion between metric space $X$ and $Y$ induces an $\epsilon$-interleaving of the persistence modules given by the d-homology of their Vietoris-Rips filtrations.
Proof. Notice that associated to a correspondence $\mathcal{C}$ between spaces $X$ and $Y$ there is a choice of a function $h:X\to Y$ whose graph is a sub-relation of $\mathcal{C}$. This map is an $\epsilon$-quasi $1$-Lipschitz map for $\epsilon=dist(\mathcal{C})$ and, hence, by Lemma 2.2 induces a map on homology as required.
In turn, we retrieve a function from $Y$ to $X$, which together with the previous gives an interleaving. $\square$
Theorem 2.7 (Stability [cohen2005stability]).
Given compact spaces $X,Y$, their interleaving distance between the associated d-homology Vietoris-Rips persistence module is bounded by twice the Gromov-Hausdorff distance between the spaces themselves.
$d_{I}(\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X),\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\leq 2\,d_{GH}(X,Y).$
This key result was first proven by Cohen-Steiner, Edelsbrunner and Harer, as duly cited, c.f. [oudot2015persistence]. It provides a central feature of persistent homology which is its stability, namely, the fact that the considered invariants are continuous on the inputs.
Let us list a few examples of naturally occurring correspondences between compact metric spaces $X,Y$ associated to an $\epsilon$-quasi $1$-Lipschitz function $f:X\to Y$.
Example 2.8.
The co-restriction of a given map between metric spaces, $f|:X\to f(X)$, being automatically surjective is already a correspondence. Its classical distortion can be interpreted as the absolute error between the distances of inputs and outputs under $f$,
$dist(f|)=sup_{x,x^{\prime}}\big|d(f(x),f(x^{\prime}))-d(x,x^{\prime})\big|.$
Example 2.9.
An inclusion of compact metric spaces $A\subseteq Y$ can be enlarged to a correspondence $C_{\iota}$ by adding all the pairs $(a,y)\in A\times Y$ for $a=argmin\penalty 10000\ d(.,y)$. Its distortion is (sharply) bounded by twice the Hausdorff distance from $A$ to $Y$, i.e.,
$dist(\mathcal{C}_{\iota})=max\big\{|d(a,a^{\prime})-d(y,y^{\prime})|:(a,y),(a^{\prime},y^{\prime})\in\mathcal{C}\big\}\leq 2\cdot d_{H}(A,Y).$
Example 2.10.
Given that a composition of correspondences is again a correspondence, the previous two examples show that a any function $f:X\to Y$ between compact metric spaces can be enlarged, as a relation, to induce a correspondence between $X$ and $Y$, so that:
$\mathcal{C}_{f}=\{(x,y)\in X\times Y\penalty 10000\ :\penalty 10000\ x=\operatorname{argmin}\penalty 10000\ d(f(.),y)\},$
so that the $x$ that pair to an $y\in Y$ are the arguments that minimize distance to $y$ via $f$.
Moreover, as a composition of the previous examples its distortion is bounded as follows:
$dist(\mathcal{C}_{f})\leq dist(f)+2\cdot d_{H}(f(X),Y).$
2.2 Bounds on the Persistent Cost
We have all the necessary ingredient to take on the subject of this work which is the persistent cost associated to a non-expanding map between compact metric spaces. We shall provide natural upper bounds that control it and stability guarantees.
Theorem 2.11.
Let $f:X\to Y$ be a 1-Lipschitz map between compact metric spaces, its persistent cost $C(f)$ on the level of the Vietoris-Rips filtration $d$-dimensional homology, satisfies the following inequality:
$C(f)\penalty 10000\ \leq\penalty 10000\ dist(f)\penalty 10000\ +\penalty 10000\ 2\cdot d_{H}(f(X),Y),$
in terms of the distortion of the original map and the Hausdorff distance between its image set and codomain.
Notice that this is a uniform bound holding for all dimensions in which the homology is computed, alike the classic stability result, Theorem 2.7.
Proof. Our argument consists of the construction a natural interleaving between $X_{\bullet}$ and $Y_{\bullet}$ associated to a $1$-Lipschitz map $f$ that in fact factors through $f_{*}$ and allows us to control kernel and cokernel modules. This is the interleaving associated to the geometric correspondence defined in Example 2.10.
We have two maps subordinated to this correspondence: the original $f\colon X\to Y$ and, in the reverse direction, a map $g\colon Y\to X$ defined as $g=f^{-1}\circ N$, where $N\colon Y\to f(X)$ picks a nearest point in the image $f(X)$ to a given $y\in Y$, and $f^{-1}$ denotes a choice of preimage under $f$. Together, they give a $(\delta+2h)$-interleaving on homology.
${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\delta+4h}(X))}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\delta+2h}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{*}}$
${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\delta+2h}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\penalty 10000\ \varphi\circ f_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{g_{*}\penalty 10000\ }$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\delta+4h}(Y))}$
In order to consider (co)-kernel modules, we recall the exact sequence (7) associated to the persistence homomorphism $f_{*}$:
$0\to\operatorname{Ker}_{\bullet}\xrightarrow{\penalty 10000\ \penalty 10000\ \penalty 10000\ }\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))\xrightarrow{\penalty 10000\ \penalty 10000\ f_{*}\penalty 10000\ \penalty 10000\ }\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\xrightarrow{\penalty 10000\ \penalty 10000\ \penalty 10000\ }\operatorname{Coker}_{\bullet}\to 0.$
One should observe that the restriction of the diagonal maps coming from the interleaving co-restrict well to the kernel.
${\operatorname{Ker}_{{}_{\bullet}}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{Ker}_{{}_{\bullet}+2\delta+4h}\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\delta+4h}(X))}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\delta+2h}(Y))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$
In doing so, we obtain an interleaving between the trivial module $\Delta$ and $\operatorname{Ker}$, using the fact that we have factored through the map $f_{*}$. Moving on, we show that $\operatorname{Coker}$ is $\delta+2h$ interleaved with the trivial module, by proving that its inner persistence map $\varphi^{\operatorname{Coker}}_{\delta+2h}$ is trivial. This follows from the previous interleaving established between $\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X))$ and $\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y))$, and the fact that this inner map is given by projection onto the cokernel, i.e., $\varphi^{\operatorname{Coker}}=\pi\circ\varphi^{Y}$. It follows that:
[…]
Where we have employed the fact that projection onto the cokernel is a persistence module homomorphism in itself. $\square$
2.3 Stability
Stability of kernel, cokernel and image persistence modules associated to a map has been claimed in [cohen2009persistent] within the general framework of sub-level set filtrations and the techniques developed in [cohen2005stability]. We give a self-contained exposition of stability for the case of Vietoris-Rips filtrations.
This motivates the definition of a Gromov-Hausdorff type distance for functions: $f_{1}:X_{1}\to Y_{1}$, $f_{2}:X_{2}\to Y_{2}$, between possibly different (compact) spaces. In order to so, we consider auxiliary $\epsilon$-isometries $S,S^{\prime},T,T^{\prime},$ between domains and codomains, respectively,as in the diagram on the side.
${X_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{1}}$$\scriptstyle{S}$${Y_{1}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T}$${X_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{f_{2}}$$\scriptstyle{S^{\prime}}$${Y_{2}\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T^{\prime}}$
Composing the given functions $f_{i}$, $i=1,2$, with the previous we may retrieve function with same domain and codomain and compute their uniform distance: $d_{\infty}\big(f_{2}\circ S,T\circ f_{1}\big)$ and $d_{\infty}\big(f_{1}\circ S^{\prime},T^{\prime}\circ f_{2}\big)$.
Definition 2.12.
For a pair of functions $f_{1}:X_{1}\to Y_{1}$, $f_{2}:X_{2}\to Y_{2}$ between compact metric spaces, consider:
$\begin{array}[]{ll}d_{GH}(f_{1},f_{2})=\underset{S,S^{\prime},T,T^{\prime}}{\inf}\max\Big\{&d_{\infty}\big(f_{2}\circ S,T\circ f_{1}\big),\penalty 10000\ \penalty 10000\ d_{\infty}\big(f_{1}\circ S^{\prime},T^{\prime}\circ f_{2}\big),\\ &\qquad\qquad dist(S),\penalty 10000\ dist(S^{\prime}),\penalty 10000\ dist(T),\penalty 10000\ dist(T^{\prime})\Big\}.\end{array}$
The infimum is taken over possible almost-isometries $S,S^{\prime},T,T^{\prime}$.
Remark 2.13.
This is a pseudo-distance between functions $f_{i}:X_{i}\to Y_{i}$ defined on possibly different compact metric spaces. Moreover,
  1. 1.
    it refines the Gromov-Hausdorff distance between respective domains and codomains,
    $2\cdot\max\{d_{GH}(X_{1},X_{2})\penalty 10000\ ,\penalty 10000\ d_{GH}(Y_{1},Y_{2})\}\leq d_{GH}(f_{1},f_{2});$
  2. 2.
    it is invariant under isometries, i.e.,
    $d_{GH}(f_{1},f_{2})=d_{GH}(f_{1}\circ S\penalty 10000\ ,\penalty 10000\ T\circ f_{2}).$
    in the case where there are true isometries $S:X_{1}\to X_{2}$, $T:Y_{1}\to Y_{2}$, between the respective (co)domains.
  3. 3.
    it induces a true distance on the class of functions related by isometries.
To the best of our knowledge this is a new distance, although it is fairly common subject on the literature to build on the Gromov-Hausdorff formalism in order to define new distances, c.f., [khezeli2023unified].
Lemma 2.14.
A pair of $1$-Lipschitz functions $f_{i}:X_{i}\to Y_{i}$, $i=1,2$, between compact metric spaces give persistent homomorphisms $f_{i}:=(f_{i})_{*}:\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X_{i}))\to\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y_{i}))$, whose induced kernel, image and cokernel persistence modules are $2\epsilon$-interleaved, provided the functions lay at distance $d_{GH}(f_{1},f_{2})\leq\epsilon$, so that:
$\max\big\{d_{I}(\operatorname{Ker}(f_{1}),\operatorname{Ker}(f_{2})),\penalty 10000\ d_{I}(\operatorname{Im}(f_{1}),\operatorname{Im}(f_{1})),\penalty 10000\ d_{I}(\operatorname{Coker}(f_{1}),\operatorname{Coker}(f_{2}))\penalty 10000\ \big\}\leq\penalty 10000\ 2\cdot d_{GH}(f_{1},f_{2})$
Proof. Fix $\epsilon$-isometries $S,S^{\prime},T,T^{\prime},$ associated to the given correspondences: $C_{X},C_{X}^{t},C_{Y},C_{Y}^{t}$, respectively, and such that $d_{\infty}\big(f_{2}\circ S,T\circ f_{1}\big),\penalty 10000\ d_{\infty}\big(f_{1}\circ S^{\prime},T^{\prime}\circ f_{2}\big)\leq\epsilon$.
Lemma 2.2 and the fact that $f_{2}\circ S$ and $T\circ f_{1}$ are both $\epsilon$-quasi Lipschitz maps at $\epsilon$ uniform distance from each other yield the following commutative diagram:
${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X_{1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{1})_{*}}$$\scriptstyle{S_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\epsilon}(X_{2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{X_{2}}_{\epsilon}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\epsilon}(X_{2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{(f_{2})_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y_{1}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{T_{*}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+\epsilon}(Y_{2}))\ignorespaces\ignorespaces\ignorespaces\ignorespaces}$$\scriptstyle{\varphi^{X_{1}}_{\epsilon}}$${\operatorname{H}_{d}(\operatorname{VR}_{{}_{\bullet}+2\epsilon}(Y_{2})).}$ (10)
In particular, $\varphi^{X_{2}}_{\epsilon}\circ S_{*}$ sends $\operatorname{Ker}((f_{1})_{*})$ into $\operatorname{Ker}((f_{2})_{*})$, since $(f_{2})_{*}$ and the corresponding $\varphi_{\epsilon}$ commute. Analogously, $\varphi^{X_{1}}_{\epsilon}\circ S^{\prime}_{*}$ sends $\operatorname{Ker}((f_{2})_{*})$ into $\operatorname{Ker}((f_{1})_{*})$. Hence, the pair $\varphi^{X_{2}}_{\epsilon}\circ S_{*}$ and $\varphi^{X_{1}}_{\epsilon}\circ(S^{\prime})_{*}$ co-restricts well to give a $2\epsilon$-interleaving between the kernels, since the original pair $S_{*},S^{\prime}_{*}$ was clearly an $\epsilon$-interleaving between the modules $\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X_{i}))$, $i=1,2$, corresponding to the domains.
Finally, a technicality. One should consider the previous argument for a sequence of $\epsilon^{\prime}>\epsilon$ converging to the true value $\epsilon=d_{GH}(f_{1},f_{2})$. Since the interleaving distance is also defined as an infimum, the lemma follows.
We omit the proof for the image and cokernel persistence modules, which make use of $\epsilon$-isometries $T,T^{\prime},$ and the inner persistence maps on the codomains. $\square$
As a corollary to the previous proof we have the following.
Corollary 2.15.
A pair of $\delta$-quasi $1$-Lipschitz functions $f_{i}:X_{i}\to Y_{i}$, $i=1,2$, between compact metric spaces give persistent homomorphisms $f_{i}:=(f_{i})_{*}:\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(X_{i}))\to\operatorname{H}_{d}(\operatorname{VR}_{\bullet}(Y_{i}))$, whose induced kernel, image and cokernel persistence modules are $2\epsilon+\delta$-interleaved, provided the functions lay at distance $d_{GH}(f_{1},f_{2})\leq\epsilon$, so that:
$\max\big\{d_{I}(\operatorname{Ker}(f_{1}),\operatorname{Ker}(f_{2})),\penalty 10000\ d_{I}(\operatorname{Im}(f_{1}),\operatorname{Im}(f_{1})),\penalty 10000\ d_{I}(\operatorname{Coker}(f_{1}),\operatorname{Coker}(f_{2}))\penalty 10000\ \big\}\leq\penalty 10000\ 2\cdot d_{GH}(f_{1},f_{2})+\delta$
Theorem 2.16.
The persistent cost $C(f)$ associated to $1$-Lipschitz functions between compact metric spaces $f,g:X\to Y$ is Lipschitz stable with respect to the invariant Gromov-Hausdorff distance between maps, more precisely, we have:
$|C(f)-C(g)|\leq 2\cdot d_{GH}(f,g).$
Proof.
$\begin{array}[]{rl}|C(f)-C(g)|&=|d_{I}(\operatorname{Ker}\oplus\operatorname{Coker}(f),\Delta)-d_{I}(\operatorname{Ker}\oplus\operatorname{Coker}(g),\Delta)|\\[8.0pt] &\leq d_{I}(\operatorname{Ker}\oplus\operatorname{Coker}(f),\operatorname{Ker}\oplus\operatorname{Coker}(g))\\[8.0pt] &\leq max\{\penalty 10000\ d_{I}(\operatorname{Ker}(f),\operatorname{Ker}(g))\penalty 10000\ ,\penalty 10000\ d_{I}(\operatorname{Coker}(f),\operatorname{Coker}(g))\penalty 10000\ \}\\[8.0pt] &\leq 2\cdot d_{GH}(f,g).\end{array}$
We have sequentially applied the triangular inequality and previous Lemma 2.14 on the final step.
$\square$
2.4 A metric example
The interleaving distance between persistence modules can be characterized via persistence homomorphisms: two persistence modules $V_{{}_{\bullet}}$ and $W_{{}_{\bullet}}$ are $\epsilon$-interleaved if and only if there exists an $\epsilon$-degree persistence homomorphism $F:V_{{}_{\bullet}}\to W_{{}_{\bullet}+\epsilon}$ with persistent cost $C(F)=2\epsilon$ ([bauer2015induced], Remark 6).
A natural question is whether such optimal homomorphisms can be realized geometrically, i.e., as maps induced by 1-Lipschitz or $\epsilon$-quasi 1-Lipschitz functions between the underlying metric spaces. The following example shows that this is not always possible, i.e., the persistent cost of any 1-Lipschitz induced map may be larger than the algebraically optimal value.
Example.
Set 1
Set 2
Figure 2: Two sets with identical barcodes throughout persistent d-homology which do not admit a $1$-Lipschitz map with small persistent cost.
Consider the finite sets presented in Figure 2 and notice that they have the same diagram in persistent homology of arbitrary dimension. In fact, the dust phase has the same 13 bars that collapse into only one at scale $\epsilon=1$ and lives on. On the $H_{1}$ level: four different classes are born at $\epsilon=1$ and die by $\epsilon=\sqrt{2}$. Higher dimensional homology is completely trivial.
However, as an exercise on non-expanding functions, one can see that it is impossible to produce such a map with trivial or at least small persistent cost in dimensions $d=0$, nor in dimension $d=1$. In fact, any $1$-Lipschitz function defined in either direction between Set 1 and Set 2 will necessarily leave one of the “corner” points unmatched and hence introduce a defect in both $H_{0}$ and $H_{1}$ with positive persistence lifespan.
Likewise, we cannot define an $\epsilon$-quasi $1$-Lipschitz map, with arbitrarily small $\epsilon>0$.